### v1

@@ Abstract
Neural image compression (NIC) has become the state-of-the-art for rate-distortion performance, yet its security vulnerabilities remain significantly less understood than those of classifiers. Existing adversarial attacks on NICs are often naive adaptations of pixel-space methods, overlooking the unique, structured nature of the compression pipeline. In this work, we propose a more advanced class of vulnerabilities by introducing T-MLA, the first targeted multiscale log–exponential attack framework. Our approach crafts adversarial perturbations in the wavelet domain by directly targeting the quality of the attacked and reconstructed images. This allows for a principled, offline attack where perturbations are strategically confined to specific wavelet subbands, maximizing distortion while ensuring perceptual stealth. Extensive evaluation across multiple state-of-the-art NIC architectures on standard image compression benchmarks reveals a large drop in reconstruction quality while the perturbations remain visually imperceptible. Our findings reveal a critical security flaw at the core of generative and content delivery pipelines.
@@ 1 Introduction
Classical image codecs, including JPEG [1], JPEG 2000 [2], BPG, and H.266/VVC [3], rely on fixed transforms and hand-crafted entropy models. However, neural image compression (NIC) [4, 5, 6, 7, 8] now matches or exceeds H.266/VVC at 0.15–0.9 bits per pixel (bpp) using standard metrics like PSNR and VIF. This progress has led to JPEG AI’s advancement to Draft International Standard, with real-time 4K decoding on mobile hardware [9]. Widespread adoption of NIC across cloud storage, content delivery networks (CDNs), and mobile applications underscores the need to understand its security vulnerabilities [10, 11, 12].
@@
[…]
Adversarial attacks on neural image compression systems are often implemented through additive perturbations in the spatial domain constrained by the $\ell_{\infty}$-norm (e.g., PGD) [19, 20], without regard to the frequency structure exploited by modern codecs [21, 22, 23, 24]. While these attacks may appear imperceptible to humans, they often inject widespread high-frequency noise that degrades compressed reconstructions.
@@
Our contributions can be summarized as follows:
  • Wavelet-Domain Multiscale Attack: We introduce a novel adversarial attack in the wavelet domain that preserves imperceptibility in both spatial and frequency representations (Figure 2). Perturbations are applied via a nonlinear log–exp mapping that adaptively modulates noise according to the local wavelet energy, preserving salient features while targeting weak or flat regions (subsection 3.3).
  • Comprehensive Evaluation: We evaluate our attack on texture-diverse images from Kodak, CLIC, and DIV2K datasets (detailed in Appendix B). Across Cheng2020-Anchor, Cheng2020-Attention, and LIC-TCM models, T-MLA consistently attains reconstruction PSNR around 24–26 dB under strict stealth constraints of 50–55 dB PSNR at the input (Figure 4(a)).
@@ 2 Related Work
Research then shifted to frequency-domain attacks [20, 29, 30], from DCT-based AdvDrop [31] to FGL [32] for mid-frequency targeting. AdvWave [33] exploited multilevel DWT for imperceptible attacks (SSIM $\approx$ 0.98), while [19] showed frequency methods improve transferability. These approaches demonstrated that targeting specific frequency bands could achieve better imperceptibility than pixel-space perturbations.
Recent benchmarks [34, 35, 36] revealed architectural robustness variations, with Cheng2020-Attention showing 15% better robustness than Cheng2020-Anchor, while JPEG AI and deeper models exhibit vulnerabilities. Various defenses combine adversarial training [27, 28, 37], wavelet-based [37, 38, 39, 40, 41] and frequency-domain approaches [42, 43, 44, 45, 46]. We outline and discuss potential defenses tailored to wavelet-domain attacks in 2.
However, while frequency-domain methods have shown promise in classification tasks, their application to neural image compression remains limited, particularly for achieving compression degradation with visual imperceptibility. For optimization-based baselines, we note that adaptive PGD variants such as APGD [47] offer stronger step-size control than classical PGD; adapting APGD to the NIC objective is a promising baseline direction, compatible with our frequency-domain design.
Scope clarification: wavelet-based attacks in other domains. Prior works have explored wavelet-domain adversarial perturbations in tasks unrelated to NIC, such as AdvWave [33] for style transfer, SITA [48] for classification attacks, and frequency-driven approaches [49] for evading similarity models. While these methods leverage frequency priors, their goals differ fundamentally from NIC’s focus on rate–distortion tradeoffs and entropy model robustness. We do not compare against these approaches, as they fall outside the NIC threat model and exploit vulnerabilities that are not present in compression pipelines. Adapting them would require redesigning their core objectives to target bitrate and reconstruction quality, which is non-trivial and misaligned with their original purpose. This clarification is important to avoid conflating adversarial objectives across domains with fundamentally different loss functions and threat assumptions.
@@ 3.1 Problem Setup
We assume a white-box setting where the attacker has full access to NIC model $f(\cdot;\theta)$. Let $\mathbf{x}\in[0,1]^{I\times J\times C}$ denote the input image and $\hat{\mathbf{x}}=f(\mathbf{x};\theta)$ the reconstructed output. The objective is to find an adversarial image $\mathbf{x}_{\mathrm{adv}}$ such that
@@
where $\mathcal{L}$ denotes a distortion loss, which can be the negative PSNR, and $\delta$ controls the perturbation magnitude. Classic spatial-domain methods, e.g., FGSM [50], PGD [51], and CW [52], optimize additive noise $\mathbf{n}$ in the spatial domain
@@ 3.2.2 Implications for Adversarial Detection
Unlike pixel-space methods, our wavelet-based attack directly manipulates frequency content. This allows fine-grained, scale- and orientation-aware perturbations that degrade compression quality while remaining visually stealthy, key for bypassing detection and surviving nonlinear NIC pipelines [53, 54].
@@ 3.3.1 Motivation for Nonlinear Perturbations
Thus, the nonlinear perturbation strategy, by scaling perturbation strength inversely with the magnitude of the original coefficients, preserves salient structures while exploiting less informative regions for noise injection. This allows for more targeted and stealthy adversarial modifications in the multiscale wavelet domain.
@@ 3.5 Adaptive Noise Bounds per Subband
Here, $\delta$ is the global noise budget, and $\alpha$ controls how the bound scales. While theory suggests $\alpha=2$ based on coefficient variance doubling at finer scales, empirical analysis on DIV2K and CLIC datasets yields $\alpha\approx 1.9$. We use $\alpha=1.8$ to balance imperceptibility and attack effectiveness (detailed analysis is provided in Appendix E).
@@ 3.6 Targeted Multiscale Log–Exp Attack (T-MLA)
Our multiscale adversarial attack is summarized in 1. At each iteration, noise is added adaptively across wavelet subbands under scale-aware constraints. The perturbations are applied in the wavelet domain using a log-exp formulation, then transformed back to pixel space via inverse DWT. The resulting adversarial image is passed through the NIC model, and the perturbation is updated to maximize a compression-targeted loss while maintaining imperceptibility.
@@ 4 Experimental Setup
Figure 3: Stealth vs. degradation tradeoff for T-MLA across NIC models and datasets. Each point represents an model-image pair and the corresponding attack performance. The axes are $\mathrm{PSNR}(x_{\mathrm{adv}},x)\uparrow$ vs. $\mathrm{PSNR}(x_{\mathrm{adv}},f(x_{\mathrm{adv}}))\downarrow$ (targets: 50/55 dB and 25 dB).
@@ 4.1 Datasets and Target Models
We evaluate on three standard image compression datasets, using representative images from each. Kodak [55] contains 24 classical $768\times 512$ PNGs with diverse textures. CLIC-pro [56] has 250 high-resolution professional photos ($\mu=2.7$MP), from which we select 63 images matched to Kodak’s local entropy distribution, then crop to $768\times 512$ (or $512\times 768$). DIV2K [57] includes 800 images at $2{,}048\times 1{,}356$; we select 65 entropy-matched samples. Selection ensures comparable complexity across datasets (see Figure B.2).
We test on three NIC models: Cheng2020-Anchor and Cheng2020-Attention [7], which use hyperprior entropy modeling (with and without attention), and LIC-TCM, a recent transformer–CNN hybrid [8].
@@ 4.2 Baseline and Evaluation Metrics
We compare our method (T-MLA) against: Pixel-Additive (PGD): a standard pixel-space PGD attack [27]; and Pixel-LogExp: our method with S=0S{=}0 and no wavelet stages, to isolate the multiscale effect.
@@ 4.3 Implementation details
PyTorch 2.4.1 (CUDA 11.8) with CompressAI library for NIC and ‘pytorch-wavelets‘ package; hardware: A100-SXM4-40GB ($\times$2) and RTX 4090 workstation. For 1 we use Haar wavelets ($S=3$ scales), Adam optimizer ($\eta=1\text{e-}2$), and perturbation budget $\delta=0.03$ with scale-adaptive constraints $\|\mathbf{N}_{k}\|_{\infty}\leq\delta_{k}$. We set $Q_{\text{in}}=55\pm 2$ dB for Cheng2020-Anchor and LIC-TCM, $Q_{\text{in}}=50\pm 2$ dB for Cheng2020-Attention, and $Q_{\text{out}}=25\pm 2$ dB for all models.
@@
   Method    Model    Stealth $\uparrow$    Attack Success $\downarrow$    PSNR    VIF    PSNR    SSIM    VIF    BPP    PGD    Cheng2020-Anchor    40.2    0.975    9.02    0.701    0.079    0.92    Cheng2020-Attention    30.7    0.817    6.44    0.587    0.066    1.51    LIC-TCM    34.4    0.920    11.3    0.728    0.138    6.71    LogExp    Cheng2020-Anchor    43.4    0.984    21.3    0.910    0.488    0.87    Cheng2020-Attention    37.0    0.968    32.0    0.944    0.929    0.93    LIC-TCM    36.6    0.946    14.7    0.825    0.223    4.40    T-MLA    Cheng2020-Anchor    55.9    0.999    24.9    0.969    0.712    0.74    Cheng2020-Attention    49.1    0.997    25.1    0.970    0.819    0.85    LIC-TCM    54.3    0.999    25.8    0.962    0.735    1.34
@@ 5.1 Quantitative Analysis
Table 1 summarizes performance across three NIC models, with results for each method-model combination averaged over the three datasets (Kodak, CLIC, DIV2K). T-MLA markedly improves stealth PSNR over pixel-space PGD while maintaining attack strength: Cheng2020-Anchor 55.89 vs. 40.15 dB ($\uparrow$15.74), Cheng2020-Attention 49.13 vs. 30.67 dB ($\uparrow$18.46), LIC-TCM 54.28 vs. 34.37 dB ($\uparrow$19.91).

### v2

@@ Abstract
Neural image compression (NIC) has become the state-of-the-art for rate-distortion performance, yet its security vulnerabilities remain significantly less understood than those of classifiers. Existing adversarial attacks on NICs are often naive adaptations of pixel-space methods, overlooking the unique, structured nature of the compression pipeline. In this work, we propose a more advanced class of vulnerabilities by introducing T-MLA, the first targeted multiscale log–exponential attack framework. We introduce adversarial perturbations in the wavelet domain that concentrate in less perceptually salient coefficients, improving the stealth of the attack. Extensive evaluation across multiple state-of-the-art NIC architectures on standard image compression benchmarks reveals a large drop in reconstruction quality while the perturbations remain visually imperceptible. On standard NIC benchmarks, T-MLA achieves targeted degradation of reconstruction quality while improving perturbation imperceptibility (higher PSNR/VIF of the perturbed inputs) compared to PGD-style baselines at comparable attack success, as summarized in our main results. Our findings reveal a critical security flaw at the core of generative and content delivery pipelines.
@@ 1 Introduction
Motivation from NIC deployment.
Classical image codecs, including JPEG [wallace1991jpeg], JPEG 2000 [christopoulos2000jpeg2000], Better Portable Graphics (BPG), and H.266/Versatile Video Coding (VVC) [bross2021vvc], rely on fixed transforms and hand-crafted entropy models. However, neural image compression (NIC) [balle2017endtoend] now matches or exceeds H.266/VVC at 0.15–0.9 bits per pixel (bpp) using standard metrics like Peak Signal-to-Noise Ratio (PSNR) and Visual Information Fidelity (VIF). This progress has led to JPEG AI’s advancement to Draft International Standard, with real-time 4K decoding on mobile hardware [jpegai2024dis]. Widespread adoption of NIC across cloud storage, content delivery networks (CDNs), and mobile applications underscores the need to understand its security vulnerabilities [Wei2021INS].
Security threats in NIC.
[…]
Throughout this work, we adopt a white-box threat model for NICs: the attacker has full gradient access to a given codec and optimizes codec-specific perturbations. Our primary goal is to provide a controlled robustness assessment of modern open-source NIC architectures, rather than to design a single universally transferable or query-efficient black-box attack across heterogeneous codecs.
@@
Limitations of existing attacks.
Adversarial attacks on neural image compression systems are often implemented through additive perturbations in the spatial domain constrained by the $\ell_{\infty}$-norm (e.g., Projected Gradient Descent (PGD)) [wang2025fba], without regard to the frequency structure exploited by modern codecs [wang2024saliency]. While these attacks may appear imperceptible to humans [liu2024freqsens], they often inject widespread high-frequency noise [Zhao2024INS] that degrades compressed reconstructions.
@@
Our specific contributions are:
  • Wavelet-Domain Multiscale Attack: We propose a novel, frequency-aware attack framework tailored for NIC. Unlike standard additive perturbations, our method employs a multiscale log–exp mapping that adapts noise energy to the local magnitude of wavelet coefficients. This design specifically targets the vulnerabilities of entropy models in NIC by hiding perturbations within the natural frequency statistics of the image, thereby achieving high stealth in both spatial and spectral domains (subsection 3.3).
  • Comprehensive Evaluation: We evaluate our attack on texture-diverse images from Kodak, CLIC, and DIV2K datasets (detailed in Appendix B). Across Cheng2020-Anchor, Cheng2020-Attention [cheng2020], and LIC-TCM [liu2023tcm] models, T-MLA consistently attains reconstruction PSNR around 24–26 dB under strict stealth constraints of 50–55 dB PSNR at the input (Figure 3).
@@ 2 Related Work
Frequency-domain adversarial attacks. Building on classical transform coding, frequency-domain attacks modify Discrete Cosine Transform (DCT) or wavelet coefficients to better exploit the internal structure of compression pipelines. Research in this area [gu2024freqdomain] has expanded from DCT-based AdvDrop [duan2021advdrop] to FGL [wang2024boosting] for mid-frequency targeting, with extensions to backdoor attacks [tan2023backdoor] and reconstruction distortion [sui2023reconattack]. AdvWave [hu2025advwave] exploited multilevel DWT for imperceptible attacks (Structural Similarity Index Measure (SSIM) $\approx$ 0.98), WaveAttack [li2024waveattack] introduced asymmetric frequency obfuscation for stealthy backdoors, while [wang2025fba] showed frequency methods improve transferability. These approaches demonstrated that targeting specific frequency bands could achieve better imperceptibility than pixel-space perturbations.
Recent benchmarks [croce2020robustbench] revealed architectural robustness variations, with Cheng2020-Attention showing 15% better robustness than Cheng2020-Anchor, while JPEG AI [kovalev2024jpegai] and deeper models exhibit vulnerabilities. Various defenses combine adversarial training [chen2021robustnic] with robustness analysis [wu2024adversarial] and gradient obfuscation [athalye2018obfuscated], wavelet-based denoising [wang2021waveletvae] and data-free methods [mustafa2024wnr], as well as frequency-domain approaches including diffusion-based defense [amerehi2025defending], contrastive learning [yang2024facl], and compression-based preprocessing [jia2019comdefend]. We outline and discuss potential defenses tailored to wavelet-domain attacks in Algorithm 2.
However, while frequency-domain methods have shown promise in classification tasks, their application to neural image compression remains limited, particularly for achieving compression degradation with visual imperceptibility. For optimization-based baselines, we note that adaptive PGD variants such as APGD [croce2020apgd] offer stronger step-size control than classical PGD; adapting APGD to the NIC objective is a promising baseline direction, compatible with our frequency-domain design.
Scope clarification: wavelet-based attacks in other domains. Prior works have explored wavelet-domain adversarial perturbations in tasks unrelated to NIC, such as AdvWave [hu2025advwave] for style transfer, SITA [kang2025sita] for classification attacks, and frequency-driven approaches [luo2022frequency] for evading similarity models. While these methods leverage frequency priors, their goals differ fundamentally from NIC’s focus on rate–distortion tradeoffs and entropy model robustness. We do not compare against these approaches, as they fall outside the NIC threat model and exploit vulnerabilities that are not present in compression pipelines. Adapting them would require redesigning their core objectives to target reconstruction quality or bitrate, which is non-trivial and misaligned with their original purpose.
@@ 3.1 Problem Setup
We assume a white-box setting where the attacker has full access to NIC model $f(\cdot;\theta)$ [balle2018hyperprior]. Let $\mathbf{x}\in[0,1]^{I\times J\times C}$ denote the input image and $\hat{\mathbf{x}}=f(\mathbf{x};\theta)$ the reconstructed output. The objective is to find an adversarial image $\mathbf{x}_{\mathrm{adv}}$ such that
@@
where $\mathcal{L}$ denotes a distortion loss, which can be the negative PSNR, and $\delta$ controls the perturbation magnitude. We deliberately do not include bitrate (BPP) in the objective: all rate values reported later are measured after optimization and serve only as descriptive diagnostics. In this setting T-MLA is optimized per codec, not as a single universal perturbation shared across architectures, which aligns with our goal of white-box robustness assessment for specific NIC models. Classic spatial-domain methods, e.g., Fast Gradient Sign Method (FGSM) [goodfellow2015explaining], PGD [madry2018towards], and Carlini & Wagner [carlini2017cw], optimize additive noise $\mathbf{n}$ in the spatial domain
@@ 3.2.2 Implications for Adversarial Detection
Unlike pixel-space methods, our wavelet-based attack directly manipulates frequency content. This allows fine-grained, scale- and orientation-aware perturbations that degrade compression quality while remaining visually stealthy, which is key for bypassing detection [liang2025frenet] and surviving nonlinear NIC pipelines [xu2025wddsr].
@@ 3.3.1 Motivation for Nonlinear Perturbations
Thus, the proposed nonlinear perturbation strategy differs from standard additive noise by scaling perturbation strength inversely with the magnitude of the original coefficients. This ensures that attacks are masked within the inherent “texture” of the image, exploiting the fact that NIC models often allocate fewer bits to high-frequency, low-energy components. By aligning the attack with these statistical properties, T-MLA achieves a targeted degradation that is structurally stealthy in the multiscale wavelet domain.
This is especially critical because modern NIC models allocate bits via entropy models conditioned on priors that expect smooth decay in high frequencies. Perturbing low-energy detail coefficients disrupts this expectation, effectively “injecting entropy” where the model expects redundancy, which forces the decoder to spend bits encoding noise or, when capacity is limited, to drop essential structural information, leading to the observed degradation.
@@ 3.5 Adaptive Noise Bounds per Subband
Here, $\delta$ is the global noise budget, and $\alpha$ controls how the bound scales. This scaling strategy applies to all subbands (including the approximation band $\mathbf{L}_{S}$), ensuring a unified optimization process that generalizes across different compression architectures without requiring model-specific subband selection. While theory suggests $\alpha=2$ based on coefficient variance doubling at finer scales, empirical analysis on DIV2K and CLIC datasets yields $\alpha\approx 1.9$. We use $\alpha=1.8$ to balance imperceptibility and attack effectiveness (detailed analysis is provided in Appendix E).
@@ 3.6 Targeted Multiscale Log–Exp Attack (T-MLA)
In this work, the term “targeted” specifically refers to driving the codec output towards a desired reconstruction PSNR level $Q_{\text{out}}$ (while maintaining a desired input PSNR level $Q_{\text{in}}$), rather than targeting a particular semantic class, spatial region, or high-level visual attribute.
Mathematically, this formulation differs from conventional distortion/perceptual losses (e.g., plain $\ell_{2}$, SSIM-, or LPIPS-based objectives) in two ways. First, the PSNR terms act as soft targets that pin the optimization to a specific operating point $(Q_{\text{in}},Q_{\text{out}})$, rather than merely encouraging “as much distortion as possible” under a global norm bound. Second, in combination with the log–exp perturbation in equation 3 and its approximation in the analysis above, the effective additive noise is reweighted by a factor $\exp(-|w|)$ in the wavelet domain, yielding a content- and scale-aware loss landscape that penalizes deviations on large-magnitude, structure-carrying coefficients much more strongly than on low-energy ones; standard $\ell_{2}$/SSIM/LPIPS losses applied in pixel space lack this built-in, coefficient-dependent weighting across scales.
Our multiscale adversarial attack is summarized in Algorithm 1. At each iteration, noise is added adaptively across wavelet subbands under scale-aware constraints. The perturbations are applied in the wavelet domain using a log-exp formulation, then transformed back to pixel space via inverse DWT. The resulting adversarial image is passed through the NIC model, and the perturbation is updated to maximize a compression-targeted loss while maintaining imperceptibility.
@@ 4 Experimental Setup
Figure 3: Stealth vs. degradation tradeoff for T-MLA across NIC models and datasets. Each point represents a model-image pair and the corresponding attack performance. The axes are $\mathrm{PSNR}(x_{\mathrm{adv}},x)\uparrow$ vs. $\mathrm{PSNR}(x_{\mathrm{adv}},f(x_{\mathrm{adv}}))\downarrow$ (targets: 50/55 dB and 25 dB).
@@ 4.1 Datasets and Target Models
We evaluate on three standard image compression datasets, using representative images from each. Kodak [franzen1999kodak] contains 24 classical $768\times 512$ PNGs with diverse textures. CLIC-pro [toderici2020clic] has 250 high-resolution professional photos, from which we select 63 images matched to Kodak’s local entropy distribution, then crop to $768\times 512$ (or $512\times 768$). DIV2K [agustsson2017ntire] includes 800 images at $2{,}048\times 1{,}356$; we select 65 entropy-matched samples. Selection ensures comparable complexity across datasets (see Figure B.2).
We test on three NIC models: Cheng2020-Anchor and Cheng2020-Attention [cheng2020], which use hyperprior entropy modeling (with and without attention), and LIC-TCM, a recent transformer–CNN hybrid [liu2023tcm].
@@ 4.2 Baseline and Evaluation Metrics
We compare our method (T-MLA) against: Pixel-Additive (PGD): a standard pixel-space PGD attack [chen2021robustnic]; and Pixel-LogExp: our method with S=0S{=}0 and no wavelet stages, to isolate the multiscale effect.
@@ 4.3 Implementation Details
PyTorch 2.7.1 (CUDA 12.6) with CompressAI library for NIC and ‘pytorch-wavelets‘ package; hardware: A100-SXM4-40GB ($\times$2) and RTX 4090 workstation. Unless otherwise stated, all experiments use a fixed random seed of 42. For Algorithm 1 we use Haar wavelets ($S=3$ scales), Adam optimizer ($\eta=1\text{e-}2$), and perturbation budget $\delta=0.03$ with scale-adaptive constraints $\|\mathbf{N}_{k}\|_{\infty}\leq\delta_{k}$. We set $Q_{\text{in}}=55\pm 2$ dB for Cheng2020-Anchor and LIC-TCM, $Q_{\text{in}}=50\pm 2$ dB for Cheng2020-Attention, and $Q_{\text{out}}=25\pm 2$ dB for all models. On an A100 GPU, a full T-MLA run for a $768{\times}512$ Kodak image takes about 20 minutes for Cheng2020-Anchor (roughly 2.5 s per 100 iterations), with comparable runtime for Cheng2020-Attention and about $5\times$ slower iterations for LIC-TCM; peak memory usage is approximately 2.0 GB (Anchor), 2.5 GB (Attention), and 11.0 GB (LIC-TCM).
@@
   Method    Model    Stealth $\uparrow$    Attack Success $\downarrow$    PSNR    VIF    PSNR    SSIM    VIF    BPP    PGD    Cheng2020-Anchor    40.15 ± 0.02    0.973 ± 0.018    8.89 ± 2.73    0.691 ± 0.103    0.081 ± 0.137    0.93 ± 0.37    Cheng2020-Attention    30.67 ± 0.04    0.815 ± 0.076    6.39 ± 3.24    0.583 ± 0.081    0.065 ± 0.181    1.52 ± 0.21    LIC-TCM    34.37 ± 0.07    0.918 ± 0.044    11.00 ± 3.17    0.717 ± 0.073    0.127 ± 0.165    6.73 ± 1.20    LogExp    Cheng2020-Anchor    43.42 ± 0.33    0.984 ± 0.004    21.32 ± 1.52    0.910 ± 0.090    0.488 ± 0.117    0.87 ± 0.35    Cheng2020-Attention    37.04 ± 0.09    0.968 ± 0.006    32.07 ± 0.78    0.944 ± 0.056    0.929 ± 0.169    0.93 ± 0.33    LIC-TCM    36.69 ± 0.75    0.946 ± 0.005    14.74 ± 1.42    0.825 ± 0.073    0.223 ± 0.106    4.40 ± 0.70    T-MLA    Cheng2020-Anchor    55.82 ± 0.87    0.999 ± 0.001    25.03 ± 0.33    0.969 ± 0.010    0.728 ± 0.140    0.85 ± 0.37    Cheng2020-Attention    49.14 ± 0.70    0.996 ± 0.003    25.20 ± 0.20    0.970 ± 0.009    0.839 ± 0.107    0.86 ± 0.37    LIC-TCM    54.31 ± 0.83    0.999 ± 0.001    25.71 ± 0.73    0.963 ± 0.011    0.735 ± 0.130    1.35 ± 0.48
@@ 5.1 Quantitative Analysis
Table 1 summarizes performance across three NIC models, with results for each method-model combination averaged over the three datasets (Kodak, CLIC, DIV2K). T-MLA markedly improves stealth PSNR over pixel-space PGD while maintaining attack strength: Cheng2020-Anchor 55.82 vs. 40.15 dB ($\uparrow$15.67), Cheng2020-Attention 49.14 vs. 30.67 dB ($\uparrow$18.47), LIC-TCM 54.31 vs. 34.37 dB ($\uparrow$19.94). For a per-codec statistical comparison on the Kodak dataset, including mean $\pm$ standard deviation and significance tests, see the bar plots in Figure I.9. For a stronger pixel-space baseline on Cheng2020-Anchor, we additionally compare against APGD [croce2020apgd] on Kodak; summary statistics are provided in Table I.4. On this setting, APGD slightly improves stealth over PGD (input PSNR $42.10$ vs. $40.15$ dB) but largely sacrifices attack strength (output PSNR increases from $8.89$ to $25.18$ dB), whereas T-MLA achieves much higher imperceptibility (PSNR $55.82$ dB, VIF $0.999$) while keeping the output PSNR around $25$ dB, yielding a strictly better stealth–success trade-off.